OpenStudy (anonymous):

Convert 17/80 into decimal

OpenStudy (anonymous):



OpenStudy (anonymous):

17/80 =0.2125

OpenStudy (anonymous):

\[\frac{ 17 }{ 80 }\times \frac{ 125 }{ 125 }=\frac{ 2125 }{ 10000 }=0.2125\]